An 18-ft board is cut into two piece, one of which is 4 ft longer that the other. HOw long is each piece?
Answer by edjones(8007)

Let the two pieces be x and x+4.
x+x+4=18
2x+4=18
2x+4-4=18-4
2x=14
2x/2=14/2
x=7'
x+4=11'
.
Ed